# 1] organization=School of Computer Sciences , addressline=Universiti Sains Malaysia, city=11800 USM Penang, country=Malaysia

### 23 Dec 2023 v5

@@ 4.1 BreakHis Dataset
The BreakHis dataset, utilized in this study focusing on breast cancer histopathological images, was generously provided by scholar Spanha [38]. This dataset comprises histopathology samples of breast cancer tumors obtained from 82 patients. The samples were sectioned under a microscope and then imaged at different magnifications (40×, 100×, 200×, 400×) to capture intricate details. Experienced medical professionals meticulously interpreted and identified the images. The dataset consists of a total of 7,909 breast cancer pathological images, comprising 2,429 benign tumor images and 5,429 malignant tumor images. Each image has dimensions of 700 × 460 pixels and is stored in PNG format with three color channels. The detailed statistics of the dataset can be found in Table 2, covering sample size, class distribution, etc. This multi-angle data resource provides a solid foundation for effective image classification and analysis research. In addition, to more intuitively display the breast cancer tissue details under different magnifications, Fig. 5 presents these pathological images as a practical example.
@@
Table 2: BreakHis dataset: benign and malignant image distribution across magnification factors
@@
Refer to caption 40x
Refer to caption 100x
Refer to caption 200x
Refer to caption 400x
Fig. 5: Images of breast cancer pathology at four magniations
4.2 Experimental environment
4.3 Experiment implementation
This research utilizes a ratio of 6:1:3 to divide the processed dataset into training, verification, and test sets, as presented in Table 3. The training set is utilized for model training, during which parameter learning occurs. The verification set is employed to assess the model’s generalization ability. Through continuous verification and parameter fine-tuning, superior model performance is ensured, with the optimal model being retained during the verification process. The test set measures the model’s recognition rate and generalization performance. Before model training, the training data is shuffled to enhance the model’s robustness and generalization capabilities.
In deep learning, the configuration of various parameters is paramount. Particularly, the learning rate plays a crucial role; an excessively high learning rate may result in the loss value exploding, while a meager learning rate may lead to overfitting and reduce the model’s convergence speed. In previous studies, the initial learning rate is usually set in the range of 0.1 to 0.0001. Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately. This is important for dealing with small and unbalanced datasets to ensure that all categories are given proper attention. In addition, the setting of the training batch size and the number of iterations is extremely important. In addition, the setting of the training batch size and the number of iterations is extremely important. Previous studies usually set the batch size between 32 and 128. Considering the GPU’s 16GB video memory limit, the lowest batch size of 32 was chosen for this study. As for the number of iterations, it is also usually set between 50 and 128. Given the small size of the BreakHis dataset, this study limited the number of iterations to 50, aiming to achieve a balance between training effectiveness and computational efficiency. Finally, this study also specifically considered the setting of the dropout rate. Previous studies generally set the dropout rate between 0.25 and 0.5. In the processing of the unbalanced BreakHis dataset, a lower dropout rate helps to ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability for these categories, thus this study chose a lower dropout rate of 0.25.
The method proposed in this study as well as the benchmark model illustrated in Table 5 apply the above settings to ensure the fairness of the experiments and the comparability of the results. These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study, the parameter settings are shown in Table 4.
Table 3: Dataset split for experimental task
@@ 4.3 Experiment implementation
Table 4: Experimental configuration
@@
4.4 Baseline selection
Table 5 presents a comparative analysis of various deep learning methods, encompassing both widely adopted approaches and the novel strategy introduced in this paper. The following section briefly outlines several of the models included in the comparison:
SEResNet50 [39]: Hu et al. proposed this model in 2018. The key feature of this model is integrating the "squeeze and excitation" attention mechanism, which aims to improve the network’s capabilities.
DenseNet-121 [40]: Huang et al. proposed this model in 2017 by introducing dense connections (dense blocks) to transfer and reuse information between network layers, thus improving the efficiency of feature transmission and reproduction. This study employs transfer learning with DenseNet-121 is employed for breast cancer classification. The advantage of using a pre-trained DenseNet is that it can adapt to small sample data.
VGG16Inception [42]: The model is presented by M. Saini and S. Susan. The article delves into a deep learning model called "VGGIN-Net," mainly applied to analyze unbalanced breast cancer datasets. The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset. To reduce training time and computational cost, the model uses a pre-trained network to migrate learned weights from one domain to another."
Table 5: Classification results of 9 network models
@@ 4.4 Baseline selection
Dual-Activated Lightweight Attention ResNet50 Model 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
@@
Dual-Activated Lightweight Attention ResNet50 Model 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
@@
Table 6: The count of parameters in the network model and the time it takes to converge.
@@
Dual-Activated Lightweight Attention ResNet50 Model 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
4.5 Experiment metrics
Used four metrics, Accuracy [46, 47, 48], F1 Score [49], IBA [50], and GMean [51], to assess the performance of the model holistically. The F1 Score is a crucial metric that combines accuracy and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances. This is especially useful for imbalanced datasets where positive and negative samples differ significantly. [52]. IBA (Imbalance Balanced Accuracy) takes a targeted approach to tackle the class imbalance challenge by harmonizing accuracy and balance. This approach offers a more holistic assessment of the model’s performance, encompassing sample diversity and fairness [53]. GMean (Geometric Mean) is a metric that calculates the geometric mean of individual classification accuracies, which is used to evaluate the model’s ability to handle category imbalance and ensure that the model can perform well across all classes [54].
In Table 5 showcases the performance of the model on the test set across various image resolutions, such as 40X, 100X, 200X, and 400X. Regardless of the solution, the model consistently outperforms nine other popular network models regarding accuracy, F1 score, IBA, and GMean. These results demonstrate that the network model has a significant advantage regardless of which image resolution is used. This gives us a deep insight into the model’s excellent performance and great potential in performing medical image classification tasks.
In Table 6, presents a comparison of the model with eight other network models regarding the number of model parameters and convergence speed. Notably, the DenseNet-121 model possesses the fewest parameters. However, the model has the shortest convergence time and the fastest average convergence speed across all image resolutions, including 40X, 100X, 200X, and 400X. This further reinforces the superior performance and potential of the model for medical image classification tasks. These results proved the efficiency and effectiveness of the model.
5 Performance evaluation and visualization
@@
5.1 Loss analysis
Fig. 6 shows the variation of training loss and validation loss with the number of training iterations for different image resolutions (including 40X, 100X, 200X, and 400X). The x-axis denotes the quantity of training iterations, while the y-axis signifies the values of the loss function. Used line plots to show the trend of training loss and validation loss and added error bars to each data point to show the standard deviation of the loss values.
@@ 5.1 Loss analysis
5.2 Accuracy analysis
@@
6 Summary and conclusions
This paper focuses on applying a lightweight attention mechanism to a ResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale datasets. The ResNet50 model incorporates the lightweight attention mechanism and a dropout layer to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper model’s performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of unbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. In future research, investigating more breast multiclassification problems is planned.
[…]

### 14 Feb 2024 v6

@@ IV-A BreakHis Dataset
The BreakHis dataset, utilized in this study focusing on breast cancer histopathological images, was generously provided by scholar Spanha [38]. This dataset comprises histopathology samples of breast cancer tumors obtained from 82 patients. The samples were sectioned under a microscope and then imaged at different magnifications (40×, 100×, 200×, 400×) to capture intricate details. Experienced medical professionals meticulously interpreted and identified the images. The dataset consists of a total of 7,909 breast cancer pathological images, comprising 2,429 benign tumor images and 5,429 malignant tumor images. Each image has dimensions of 700 × 460 pixels and is stored in PNG format with three color channels. The detailed dataset statistics in Table 3 cover sample size, class distribution, etc. This multi-angle data resource provides a solid foundation for effective image classification and analysis research. In addition, to more intuitively display the breast cancer tissue details under different magnifications, Fig. 6 presents these pathological images as a practical example.
Refer to caption 40x
Refer to caption 100x
Refer to caption 200x
Refer to caption 400x
Fig. 6: Images of breast cancer pathology at four magniations
@@
TABLE III: BreakHis dataset: benign and malignant image distribution across magnification factors
@@
TABLE IV: Dataset split for experimental task
@@
TABLE V: Experimental configuration
@@
IV-B Experimental environment
IV-C Experiment implementation
This research utilizes a ratio of 6:1:3 to divide the processed dataset into training, verification, and test sets, as presented in Table 4. The training set is used for model training, during which parameter learning occurs. The verification set is employed to assess the model’s generalization ability. Continuous verification and parameter fine-tuning ensure superior model performance, with the optimal model pre-trained during verification. The test set measures the model’s recognition rate and generalization performance. Before model training, the training data is shuffled to enhance the model’s robustness and generalization capabilities.
In deep learning, the configuration of various parameters is paramount. Notably, the learning rate plays a crucial role; an excessively high learning rate may result in the loss value exploding, while a meager learning rate may lead to overfitting and reduce the model’s convergence speed. In previous studies, the initial learning rate is usually set in the range of 0.1 to 0.0001. Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately. This is important for dealing with small and unbalanced datasets to ensure that all categories are given proper attention. In addition, setting the training batch size and the number of iterations is significant. In addition, setting the training batch size and the number of iterations is substantial. Previous studies usually set the batch size between 32 and 128. Considering the GPU’s 16GB video memory limit, the lowest batch size of 32 was chosen for this study. As for the number of iterations, it is also usually set between 50 and 128. Given the size of the BreakHis dataset, this study limited the number of iterations to 50, aiming to balance training effectiveness and computational efficiency. Finally, this study also specifically considered the dropout rate setting. Previous studies generally set the dropout rate between 0.25 and 0.5. In processing the unbalanced BreakHis dataset, a lower dropout rate helps ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability. Therefore, this study chose a lower dropout rate of 0.25.
The method proposed in this study, and the benchmark model illustrated in the Table apply the above settings to ensure the experiments’ fairness and the comparability of the results. These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study; the parameter settings are shown in Table 5.
TABLE VI: Classification results of 9 network models
@@ IV-C Experiment implementation
DALAResNet50 Model 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
@@
DALAResNet50 Model 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
@@
TABLE VII: The count of parameters in the network model and the time it takes to converge.
@@
DALAResNet50 Model 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
Note: In the table, the DenseNet-121 model indicates that a pre-trained DenseNet-121 model was employed for transfer learning.
IV-D Baseline selection
Table 6 presents a comparative analysis of various deep learning methods, encompassing widely adopted approaches and the novel strategy introduced in this paper. The following section briefly outlines several of the models included in the comparison:
SEResNet50 [39]: Hu et al. proposed this model in 2018. The key feature of this model is integrating the ”squeeze and excitation” attention mechanism, which aims to improve the network’s capabilities.
DenseNet-121 [40]: Huang et al. proposed this model in 2017 by introducing dense connections (dense blocks) to transfer and reuse information between network layers, thus improving the efficiency of feature transmission and reproduction. This study employs transfer learning with DenseNet-121, which is utilized for classifying breast cancer. The advantage of using a pre-trained DenseNet is that it can adapt to small sample data.
VGG16Inception [42]: The model is presented by M. Saini and S. Susan. The article delves into a deep learning model called ”VGGIN-Net,” which is mainly applied to analyze unbalanced breast cancer datasets. The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset. To reduce training time and computational cost, the model uses a pre-trained network to migrate learned weights from one domain to another.”
IV-E Experiment metrics
Used four metrics, Accuracy [46, 47, 48], F1 Score [49], IBA [50], and GMean [51], to assess the performance of the model holistically. The F1 Score is a crucial metric that combines accuracy and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances. This is especially useful for imbalanced datasets where positive and negative samples differ significantly [52]. IBA (Imbalance Balanced Accuracy) takes a targeted approach to tackle the class imbalance challenge by harmonizing accuracy and balance. This approach offers a more holistic assessment of the model’s performance, encompassing sample diversity and fairness [53]. GMean (Geometric Mean) is a metric that calculates the geometric mean of individual classification accuracies, which is used to evaluate the model’s ability to handle category imbalance and ensure that the model can perform well across all classes [54].
Table 6 showcases the performance of the model on the test is set across various image resolutions, such as 40X, 100X, 200X, and 400X. Regardless of the solution, the model consistently outperforms nine other popular network models regarding accuracy, F1 score, IBA, and GMean. These results demonstrate that the network model has a significant advantage regardless of which image resolution is used. This gives us a deep insight into the model’s excellent performance and great potential in performing medical image classification tasks.
Table 7 compares the model with eight other network models regarding the number of model parameters and convergence speed. Notably, the DenseNet-121 model possesses the fewest parameters. However, the model has the shortest convergence time and the fastest average convergence speed across all image resolutions, including 40X, 100X, 200X, and 400X. This further reinforces the superior performance and potential of the model for medical image classification tasks. These results proved the efficiency and effectiveness of the model.
V Performance evaluation and visualization
@@
V-A Loss analysis
Fig. 7 shows the variation of training loss and validation loss with the number of training iterations for different image resolutions (including 40X, 100X, 200X, and 400X). The x-axis denotes the quantity of training iterations, while the y-axis signifies the values of the loss function. Line plots were used to show the trend of training loss and validation loss, and error bars for each data point were used to indicate the standard deviation of the loss values.
V-B Accuracy analysis
@@
VI Summary and conclusions
This paper focuses on applying a lightweight attention mechanism to a ResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale datasets. The ResNet50 model incorporates the lightweight attention mechanism and a dropout layer to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper model’s performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of unbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. In future research, investigating more investigations of breast multiclassification problems are planned.